A person has near point at 100 cm. What power of lens is needed to read at 20 cm if he/she uses (a) contact lens, (b) spectacles having glasses 2.0 cm separated from the eyes?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. The person has near point 100 cm. It is needed to read at a distance of 20cm. (a) When contact lens is used, u = – 20 cm = – 0.2m, v = – 100 cm = –1 m So, 1/f = 1/v – 1/u = 1/-1 – 1/-0.2 = - 1 + 5 = + 4D (b) When spectacles are used, u = – (20 – 2) = – 18 cm = – 0.18m, v = – 100 cm = –1 m So, 1/f = 1/v – 1/u = 1/-1 – 1/-0.18 = - 1+ 5.55 = + 4.5D